# Yanjie Pan1∗    Qingdong He2∗    Zhengkai Jiang3    Pengcheng Xu4    Chaoyi Wang5    Jinlong Peng2 Haoxuan Wang1    Yun Cao2    Zhenye Gan2    Mingmin Chi1    Bo Peng1    Yabiao Wang2

### v1

@@ 4.1 Main Results
Comparison of CLIP Score. To investigate the impact of additional visual controls on text adherence, Tab. 1 compares CLIP Score metric across methods under maximal visual conditional controls. The results reveal a consistent trade-off: methods with fewer visual constraints achieve higher text adherence. Specifically, PixelPonder demonstrates superior performance under full conditional control, achieving CLIP scores of 78.60 (MultiGen-20M) and 77.12 (Subject-200K), outperforming the compared methods. Although Cocktail attains higher scores (79.87 and 79.38, respectively) under significantly fewer visual constraints, the strong text adherence of Cocktail significantly compromises the consistency of its visual performance. As illustrated in Fig. 5, the images generated by Cocktail exhibit pronounced issues of object loss and mutations when compared to those produced by other methods. This highlights PixelPonder’s unique capability to balance precise visual control with competitive text adherence, even under the most challenging full-condition settings.
Condition Types Comparison. To examine the effect of the number of visual control types on generation, Fig. 6 shows the visual results under various condition combinations, with qualitative evaluation metrics provided in Tab. 2. The examples in Fig. 6 highlight that single-condition control methods exhibit inconsistencies and instability in the generated images, such as the background lotus leaves in row three, which display varying layouts and shapes under Canny, depth, and HED control. In contrast, PixelPonder maintains high consistency in visual features across all condition combinations. Furthermore, in Tab. 2, ControlNet++ shows significant fluctuations in FID and SSIM scores across the three visual conditions. In comparison, PixelPonder demonstrates minimal variation in these metrics. Notably, PixelPonder outperforms the best single-condition generation in both SSIM and FID under full-condition control, while maintaining strong performance on the MUSIQ metric.
4.2 Ablation Study
Patch Selection. In Fig. 7, we illustrate the statistical distribution of each visual condition patch during the denoising steps. During the initial denoising stages from step 0 to 5, the denoising process comes from low-frequency regions to high-frequency regions. However, from step 5 to 25, the trend is completely opposite. This suggests that after establishing a rough layout in the initial stages, generating low-frequency or high-frequency regions does not necessarily require consistent information with the layout. On the contrary, it requires visual information from other frequencies to leave sufficient space to meet other visual requirements. More discussion is shown in the supplementary material.
Patch Size. To address the issue of overlapping in visual control signals, we propose a patch-level visual condition adaptation method. Tab. 3 presents the performance of PixelPonder under different patch sizes. The experimental results indicate that finer-grained image condition combination can better resolve conflicts among various visual elements in the visual conditions.
@@ 4.2 Ablation Study
5 Conclusion
In this paper, we present PixelPonder, a novel framework for compositional visual conditioning in diffusion-based image generation. Our key contribution lies in addressing the fundamental limitation of existing methods in handling conflicting guidance from multiple control signals. Specifically, we introduce two novel components, a Patch-level Adaptive Condition Adaptation mechanism that dynamically resolves spatial conflicts through learnable attention gates, and a Time-aware Control Injection scheme that orchestrates condition influence across denoising phases. Extensive experiments on several benchmarks demonstrate significant improvements over state-of-the-art methods. With PixelPonder, users can depict different aspects of objects by utilizing various visual conditions, thereby accurately realizing their various creations.
@@ 5 Conclusion
The supplementary material presents the following sections to strengthen the main manuscript:
  • Implementation details.
  • More experiments about scale study.
  • Patch Selection Analysis
  • Generation Stability Comparison
  • Image Consistency Exploration
  • More qualitative results.
A1 Implementation Details
Dataset. Our training dataset is randomly sampled from the MultiGen-20M [20] dataset, resulting in a training set of 2.5 million samples. Each training sample contains an input image $x$, a text prompt, and a set of visual conditions $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{n}\}$. This set $\mathcal{C}$ of visual conditions includes various types of visual cues, such as canny maps, sketches, depth maps, and pose maps, obtained from several other methods [47, 31, 3, 45]. In the testing dataset, due to the small size of the MultiGen-20M test set, and to ensure fairness in comparison, we combine the test and validation sets of MultiGen-20M to form a test set of up to 5,500 text-image pairs. Additionally, to increase the confidence in our conclusions, we randomly sample 5,000 text-image pairs from the Subject-200K[40] dataset as a second test set. This dataset provides more complex text compared to MultiGen-20M, and the Subject-200Ks in each image are clearly defined.
Training Setting. We conduct the training on 24 Nvidia V100 GPUs, using a batch size of 3 for every 3 GPUs, for a total of 200,000 training steps. The AdamW optimizer is employed with a learning rate of $2\times 10^{-5}$. The default setting for patch size is 2, and the default number of patches selected per $U_{i}$ is set to $\frac{W}{\text{patch size}}$ (where $W$ is the width of the image features). The default depth of the ISB module is 1, and the default number of Zero DSBs and Zero SSBs is set to 2 and 4, respectively. During training, we randomly select 3 out of 4 types of conditions (canny maps, depth maps, sketch maps, and pose maps) as visual control inputs. Regarding the pose condition, since not all images contain humans, we do not discard the pose condition to ensure its effective training.
@@
Evaluation Metrics. For multi-visual condition control tasks, we evaluate generation quality, controllability, and text-image consistency using a comprehensive set of mathematical and model-based metrics. Generation quality is assessed through FID[14] and MUSIQ [18]. FID, a model-based metric, quantifies the similarity between feature distributions of real and generated images, providing a robust measure of image fidelity. MUSIQ, another perceptual model-based metric, evaluates image quality from a human perceptual perspective by capturing both local and global quality attributes. Controllability is measured using SSIM [43], a mathematically grounded metric that computes the structural similarity between reference and generated images, offering a precise assessment of visual fidelity. Finally, text-image consistency is evaluated via CLIP Score [13, 28], a model-based metric that measures the semantic alignment between input text and generated images.
A2 Scale Study
@@
Data Scale. We evaluate the performance of PixelPonder on different scales of training datasets, with the results shown in Tab. A1. Due to limitations in computational resources and time, we only explored up to 2.5 million data. As shown in the table, PixelPonder demonstrates significant improvements in FID and SSIM as the scale of the training dataset increases. However, MUSIQ, constrained by the original reference images, does not show further improvement after a substantial increase in the size of the data set. Compared to other methods, such as ControlNet++ and UniControl, our method achieves results comparable to those obtained by tens of millions of training samples, while being trained on only millions of samples. This indicates that there is still room for improvement in the precise control capabilities of PixelPonder.
@@
Table A1: Ablation on data scale. FID, SSIM and MUSIQ are reported on MultiGen-20M datasets. $k$ represents thousand, and $m$ represents million.
A3 Patch Selection Analysis
@@
Table A2: Ablation on fourier. × represents the non-use of fourier, while the ✓ represents the use of fourier.
@@
where $t$ is the current time step, $M_{low}$ is a low-frequency mask, $\tilde{M}_{low}$ is a high-frequency mask, $\alpha$ is the correction hyperparameter, $f$ is the spectrogram, and $f_{\text{low}}$ and $f_{\text{high}}$ represent the corrected low-frequency and high-frequency signals, respectively. $f_{\text{low}}$ decreases over time steps, while $f_{\text{high}}$ increases. We incorporate this Fourier correction into the control network, and the results are shown in Tab. A2. It demonstrates that the image quality generated by the Fourier correction constructed under the previous conclusions has declined in various aspects, which indirectly supports the validity of our conclusions.
@@
In the two examples shown in Fig. 4, due to the fine-grained division of patches, the pose maps are mostly blank noise at many patch locations, resulting in a lack of control effect. However, contrary to our expectations, in the invalid areas corresponding to the pose maps, patch selection still chooses the blank noise from the pose maps to perturb the generation, which leads to better generation results. This is supported by the comparisons in Tab. 2, where the generation results under all conditional control scenarios significantly outperform other combinations of conditions, particularly in consistency evaluation metrics such as FID and SSIM. This indicates that appropriately adding noise during the denoising process does not negatively impact image generation; rather, it frees up more space for other visual condition controls to exert their influence, which is consistent with our conclusions regarding the selection tendencies of various visual conditions in patch selection.
A4 Generation Stability Comparison
@@
Figure A1: Comparison of multiple random generations under single-condition and multi-condition. Ours employs all available conditions, whereas canny and depth denote the use of only the canny or depth conditions, respectively.
@@
As shown in Fig. A1, the previous single-condition control framework encounters significant challenges in preserving both stability and consistency regarding the detailed features and morphological characteristics of target objects across multiple generated images. Particularly in the cake and furniture scenarios, Uni-ControlNet manifests several limitations, including texture degradation, contour deformation, and background inconsistency. In stark contrast, PixelPonder demonstrates superior performance in maintaining both consistency and stability across these dimensions. A notable example can be observed in the cake scenario, where PixelPonder effectively maintains precise control over the cream’s morphological features while simultaneously ensuring consistent dimensional attributes of the cake across all generated images.
A5 Image Consistency Exploration
@@
To investigate whether the images generated by PixelPonder are consistent with the visual control conditions in terms of visual elements, we provide additional relevant visual demonstrations in Fig. A2 and Fig. A3.
@@
Figure A2: Image consistency comparison between official or re-implemented methods and our proposed model.
Figure A3: Image consistency comparison between official or re-implemented methods and our proposed model.
A6 More Qualitative Comparison
In Fig. A4 and Fig. A5, we present more visual results across various scenes. Previous methods do not separate multiple classes of visual conditions in visual control; instead, they inject multiple visual conditions simultaneously at a single time step, redundantly introducing visual elements of cannys. Our proposed PixelPonder, through patch-level separation of visual conditions and temporal domain separation, avoids the generation of overlapping conflicting visual signals and achieves the injection of complementary information among various visual elements in the temporal domain, resulting in superior generation performance under multi-visual condition control.
Figure A4: Visualization comparison between official or re-implemented methods and our proposed model in different conditional controls.
Figure A5: Visualization comparison between official or re-implemented methods and our proposed model in different conditional controls.

### v2

@@ Introduction
Ablation Study
Patch Selection. In Fig.7, we illustrate the statistical distribution of each visual condition patch during the denoising steps, based on tens of millions of samples. The range across all stages exceeds 1%, indicating a significant difference between the Patch Selection strategy and random sampling. Furthermore, The trends show that generating low-frequency or high-frequency regions does not require consistency with the initial layout; instead, it benefits from incorporating visual information from other frequency domains to meet diverse visual requirements. Visual comparisons are provided in the supplementary material.
@@
Patch Size. Tab.3 presents the performance of PixelPonder under different patch sizes. The experimental results indicate that finer-grained image condition combination can better resolve conflicts among various visual elements in the visual conditions.
@@
Conclusion
In this paper, we present PixelPonder, a novel framework for compositional visual conditioning in diffusion-based image generation. Our key contribution lies in addressing the fundamental limitation of existing methods in handling redundant guidance from multiple controls. Specifically, we introduce two novel components, a Patch-level Adaptive Condition Adaptation mechanism that dynamically resolves spatial conflicts through learnable attention gates, and a Time-aware Control Injection scheme that orchestrates condition influence across denoising phases. Extensive experiments on several benchmarks demonstrate significant improvements over state-of-the-art methods. With PixelPonder, users can depict different aspects of objects by utilizing various visual conditions, thereby accurately realizing their creations.
@@ Conclusion
Implementation Details
Dataset. Our training dataset is randomly sampled from the MultiGen-20M (Li et al. 2025) dataset, resulting in a training set of 2.5 million samples. Each training sample contains an input image $x$, a text prompt, and a set of visual conditions $\mathcal{C}=\{C_{1},C_{2},\ldots,C_{n}\}$. This set $\mathcal{C}$ of visual conditions includes various types of visual cues, such as canny maps, sketches, depth maps, and pose maps, obtained from several other methods (Yang et al. 2023b; Ranftl et al. 2020; Canny 1986; Xie and Tu 2015). In the testing dataset, due to the small size of the MultiGen-20M test set, and to ensure fairness in comparison, we combine the test and validation sets of MultiGen-20M to form a test set of up to 5,500 text-image pairs. Additionally, to increase the confidence in our conclusions, we randomly sample 5,000 text-image pairs from the Subject-200K(Tan et al. 2024) dataset as a second test set. This dataset provides more complex text compared to MultiGen-20M, and the Subject-200Ks in each image are clearly defined.
Training Setting. We conduct the training on 8 GPUs with 96GB of VRAM, using a batch size of 4 for every GPU, for a total of 200,000 training steps. The AdamW optimizer is employed with a learning rate of $2\times 10^{-5}$. The default setting for patch size is 2, and the default number of patches selected per $U_{i}$ is set to $\frac{W}{\text{patch size}}$ (where $W$ is the width of the image features). The default depth of the ISB module is 1, and the default number of Zero DSBs and Zero SSBs is set to 2 and 4, respectively. During training, we randomly select 3 out of 4 types of conditions (canny maps, depth maps, sketch maps, and pose maps) as visual control inputs. Regarding the pose condition, since not all images contain humans, we do not discard the pose condition to ensure its effective training.
@@
Evaluation Metrics. For multi-visual condition control tasks, we evaluate generation quality, controllability, and text-image consistency using a comprehensive set of mathematical and model-based metrics. Generation quality is assessed through FID(Heusel et al. 2017) and MUSIQ (Ke et al. 2021). FID, a model-based metric, quantifies the similarity between feature distributions of real and generated images, providing a robust measure of image fidelity. MUSIQ, another perceptual model-based metric, evaluates image quality from a human perceptual perspective by capturing both local and global quality attributes. Controllability is measured using SSIM (Wang et al. 2004), a mathematically grounded metric that computes the structural similarity between reference and generated images, offering a precise assessment of visual fidelity. The evaluation metric for inference speed is the time cost per image, measured in seconds per image (s/per image). Finally, text-image consistency is evaluated via CLIP Score (Hessel et al. 2021; Radford et al. 2021), a model-based metric that measures the semantic alignment between input text and generated images.
Scale Study
@@
Data Scale. We evaluate the performance of PixelPonder on different scales of training datasets, with the results shown in Tab.5. Due to limitations in computational resources and time, we only explored up to 2.5 million data. As shown in the table, PixelPonder demonstrates significant improvements in FID and SSIM as the scale of the training dataset increases. However, MUSIQ, constrained by the original reference images, does not show further improvement after a substantial increase in the size of the data set. Compared to other methods, such as ControlNet++ and UniControl, our method achieves results comparable to those obtained by tens of millions of training samples, while being trained on only millions of samples. This indicates that there is still room for improvement in the precise control capabilities of PixelPonder.
@@
Table 5: Ablation on data scale. FID, SSIM and MUSIQ are reported on MultiGen-20M datasets. $k$ represents thousand, and $m$ represents million.
Patch Selection Analysis
@@
Table 6: Ablation on fourier. × represents the non-use of fourier, while the ✓ represents the use of fourier.
@@
where $t$ is the current time step, $M_{low}$ is a low-frequency mask, $\tilde{M}_{low}$ is a high-frequency mask, $\alpha$ is the correction hyperparameter, $f$ is the spectrogram, and $f_{\text{low}}$ and $f_{\text{high}}$ represent the corrected low-frequency and high-frequency signals, respectively. $f_{\text{low}}$ decreases over time steps, while $f_{\text{high}}$ increases. We incorporate this Fourier correction into the control network, and the results are shown in Tab.6. It demonstrates that the image quality generated by the Fourier correction constructed under the previous conclusions has declined in various aspects, which indirectly supports the validity of our conclusions.
@@
In the examples shown in Fig.4 of the main paper, due to the fine-grained division of patches, the pose maps are mostly blank noise at many patch locations, resulting in a lack of control effect. However, contrary to our expectations, in the invalid areas corresponding to the pose maps, patch selection still chooses the blank noise from the pose maps to perturb the generation, which leads to better generation results. This is supported by the comparisons in Tab.2 of the main paper, where the generation results under all conditional control scenarios significantly outperform other combinations of conditions, particularly in consistency evaluation metrics such as FID and SSIM. This indicates that appropriately adding noise during the denoising process does not negatively impact image generation; rather, it frees up more space for other visual condition controls to exert their influence, which is consistent with our conclusions regarding the selection tendencies of various visual conditions in patch selection.
Generation Stability Comparison
@@
Figure 8: Comparison of multiple random generations under single-condition and multi-condition. Ours employs all available conditions, whereas canny and depth denote the use of only the canny or depth conditions, respectively.
Figure 9: Generation quality comparison between Random Selection and Patch Adaptation. Random selection can lead to a loss of information, resulting in reduced capability to generate details and clarity.
@@
As shown in Fig.8, the previous single-condition control framework encounters significant challenges in preserving both stability and consistency regarding the detailed features and morphological characteristics of target objects across multiple generated images. Particularly in the cake and furniture scenarios, Uni-ControlNet manifests several limitations, including texture degradation, contour deformation, and background inconsistency. In stark contrast, PixelPonder demonstrates superior performance in maintaining both consistency and stability across these dimensions. A notable example can be observed in the cake scenario, where PixelPonder effectively maintains precise control over the cream’s morphological features while simultaneously ensuring consistent dimensional attributes of the cake across all generated images.
As shown in Fig.9, randomly selected visual results exhibit significant deviations in terms of image layout, fine details, and intra-image contextual understanding of the garment-person interaction. Although the statistical difference is only a few percentage points, the visual discrepancy is highly noticeable. This demonstrates the effectiveness of the patch adaptation strategy in the generation process.
Image Consistency Exploration
@@
To investigate whether the images generated by PixelPonder are consistent with the visual control conditions in terms of visual elements, we provide additional relevant visual demonstrations in Fig.10 and Fig.11.
@@
Figure 10: Image consistency comparison between official or re-implemented methods and our proposed model.
Figure 11: Image consistency comparison between official or re-implemented methods and our proposed model.
More Qualitative Comparison
In Fig.10 and Fig.11, we present more visual results across various scenes. Previous methods do not separate multiple classes of visual conditions in visual control; instead, they inject multiple visual conditions simultaneously at a single time step, redundantly introducing visual elements of cannys. Our proposed PixelPonder, through patch-level separation of visual conditions and temporal domain separation, avoids the generation of overlapping conflicting visual signals and achieves the injection of complementary information among various visual elements in the temporal domain, resulting in superior generation performance under multi-visual condition control.